The profit made by a company when 60 units of its product is sold is R 1 600.00. When 150 units of its products are sold, the profit increases to R 5 200.00. Assuming that the profit function is linear and of the form
1
Expert's answer
2019-04-02T11:11:45-0400
"P(u)=a+bu"

"P(60)=a+b(60)=R\ 1 600.00"

"P(150)=a+b(150)=R\ 5 200.00"

"90b=R\ 3600.00 \Rarr b=R\ 40.00"

"a=R\ 1 600.00-R\ 40.00(60)=-R\ 800.00"

"P(u)=-800+40u"

Break-even level: total profit is zero


"P(u)=0 \Rarr -800+40u=0, u=20"

Number of units that need to be sold to realise a profit of R 12 000.00.


"-800+40u=R\ 12 000.00"

"u=280"